Nandu starts journey on boat. When wind blows his cap falls and starts floating downstream in opposite direction of boat (before nandu finds out). Boat continuously going upstream for 12 min. and comes in downstream. When cap reaches starting point, nandu catches it. If cap of nandu has fallen 3 km away from starting point, find speed of river flow?


1
Expert's answer
2021-02-24T07:07:53-0500


Let

The speed of boat is - A

The speed of current - B

YZ distance= boat upstream time × boat upstream speed

=12(A-B)

ZX distance = t(A+B)

XY distance = (12+t)B

XY+YZ=ZX

(12+t)B+12(A-B)=t(A+B)

12B+tB+12A-12B=tA+tB

12B cancelled by -12B and tB cancelled by tB , after this cancellation we get 12A=tA ,

So we get t= 12.

t=12 , t is time take by Nandu

Total time take by Nandu is 12+12=24 min so t=24


Speed of river flow "Speed= \frac {distance}{time}
 ,S=\frac {3}{24
}"Kmpm

"S= \frac {3}{24}×60"

S= 7.5 kmph